OpenStudy (anonymous):

I'm trying to write a paper for my physics class, and this is what my teacher is asking me to include:

OpenStudy (anonymous):

Derive the work-energy theorem starting with \[a = F/m\] and \[Vf^2 = Vi^2 +2ad\] and ending with \[W=DeltaKE\]

OpenStudy (danjs):

Imagine throwing a ball vertically in a gravitational field. The initial velocity is Vi and the final velocity is Vf and during that time interval t, the ball travels a distance d. Using the kinematic equation you provide to describe the motion, the acceleration, a, is g. Using Newton's second law you provided, F=ma...The only force here is gravity F=-mg If you recall from earlier topics, you conclude from kinematics that the height the ball reaches is found to be: height = d = Vi^2/(2*g) -----rearranging to get Vi^2/2 = g*d -----multiply both sides by m 1/2(m*Vi^2) = m*g*d recall: Work = Force *distance = (-)m*g*d KE = 1/2(m*V^2) 1/2(m*Vi^2) = m*g*d Work = KE In this example, the height the ball travels in the gravity field is d, and when the ball reaches the top, the KE is zero. Work = Delta(KE)

OpenStudy (danjs):

\[Wa \rightarrow b =\int\limits_{a}^{b} F*dx\] \[F=m*a=m*\frac{ dv }{ dt } \] \[dx = v*dt\] \[Wa \rightarrow b = \int\limits_{a}^{b}m*\frac{ dv }{ dt }*v*dt = \int\limits_{a}^{b}m*v*dv \] \[Wa \rightarrow b = m*1/2*Vb ^{2} - m*1/2*Va ^{2} \]